OpenStudy (anonymous):

5/3d=15/9 please help! soldve the equation. use the domain of rational numbers.

OpenStudy (anonymous):



OpenStudy (anonymous):

We can cross multiply to find: \[ 5\cdot9=15\cdot3d=45d=45 \]So: \[ d=1 \]Another way would be: \[ \frac{15}{9}=\frac{5\cdot\not3}{3\cdot\not3}=\frac{5}{3} \]So, \(d\) would have to be one.

OpenStudy (anonymous):

uhm d=1?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ooo got it thanks